Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability question

Expert replies
by MTariq » Tue Mar 16, 2010 9:24 am
Can someone point out the mistake. what am I doing wronge? Answer should be the same using both approachesA machine is made up of two components, A and B. Each component either works or fails. The failure or non-failure of one component is independent of the failure or non-failure of the other component. The machine works if at least one of the components works. If the probability that each component works is 2/3, what is the probability that the machine works?

Prob that neither component works = 1/3*1/3
Prob that machine works = 1- 1/9 = 8/9

Prob Machine works = Prob (A works) * Prob (B does not work) + Prob (B works) * Prob (A does not work)
= 2/3*1/3 + 2/3*1/3
= 2/9 + 2/9 = 4/9
Join the discussion
Source: — Problem Solving |

by kstv » Tue Mar 16, 2010 9:48 am
So prob of the machine working is when at least one of the comp. works. But what about the probability that both comp. might work. U are not accounting for it in the second method.

while it is being assumed in your calculation in the first method. Probability of both working is 2/3*2/3 = 4/9 the diff in the two methods.


think of components as switches it is either on or off.
Join the discussion

by MTariq » Tue Mar 16, 2010 10:11 am
Ohhkay, I was missing out the third scenerio.
Got it, Thanks!
Join the discussion

by hooliganpete » Fri Mar 19, 2010 11:14 am
The reason above seems incorrect to me but maybe someone can elaborate. If the machine can function with only one component functioning, then the only condition that would render the machine inoperable would be when neither components A nor B work.

The probability of either not working is 1/3. Thus, the probability that neither works is (1/3)(1/3)=1/9. Probability of the machine functioning minus the only condition which makes it not function is 1-(1/9)=8/9.

Can someone explain why, if at all, the above reasoning is incorrect?
Join the discussion

by harshavardhanc » Fri Mar 19, 2010 12:11 pm
hooliganpete wrote:The reason above seems incorrect to me but maybe someone can elaborate. If the machine can function with only one component functioning, then the only condition that would render the machine inoperable would be when neither components A nor B work.

The probability of either not working is 1/3. Thus, the probability that neither works is (1/3)(1/3)=1/9. Probability of the machine functioning minus the only condition which makes it not function is 1-(1/9)=8/9.

Can someone explain why, if at all, the above reasoning is incorrect?

Pete,

Your reasoning is the first method employed by MTariq, which is one of the correct ways to reach the answer.

But, he was not getting the same answer in the second method, as he didn't consider the case where both the machines work. That's what kstv explained in his post. That probability will be 4/9, which when added to the two cases will fetch the final answer.
Regards,
Harsha
Join the discussion

by hooliganpete » Fri Mar 19, 2010 1:28 pm
Thanks Harsha... as long as I'm getting the correct answer I suppose I can forgive myself for not paying close enough attention!
Join the discussion